# Roman Garipov $^{*\,\dagger}$

### v3

@@ 1 Introduction
Recent advances in LLM capabilities, including chain-of-thought reasoning (Wei et al., 2022; Kojima et al., 2022; Suzgun et al., 2022), writing complex software (Rozière et al., 2023; Li et al., 2023; Jiang et al., 2024), or interacting with external tools (Schick et al., 2023; Qin et al., 2023), increasingly rely on inference-time computation (Snell et al., 2024; Beeching et al., 2024). This progress is further accelerated with the release of reasoning-capable models, both proprietary (OpenAI et al., 2024; Anthropic, 2024; Google DeepMind, 2025) and open-access (DeepSeek-AI et al., 2025; Meta, 2025; Qwen Team, 2025), that were explicitly trained to perform these kinds of inference-time computation. However, as LLMs tackle harder problems, they also tend to generate longer sequences (Muennighoff et al., 2025) with tens of thousands of tokens (Yeo et al., 2025), taking up tens of minutes (and hundreds of dollars in costs) per task (ARC Prize Foundation, 2024).
A popular way to speed up LLM inference is through speculative decoding (Leviathan et al., 2023; Chen et al., 2023), which uses a small “draft” model to propose the likely next tokens, then verifies these tokens with the main model in parallel. This method, along with its successors (Miao et al., 2023; Cai et al., 2024; Li et al., 2024d), can speed up LLM inference while guaranteeing that the generated outputs match those of the original model (for greedy inference) or follow the same distribution. To achieve this, speculative decoding algorithms check if the draft tokens match the original model predictions. If there is a mismatch, they discard the incorrect token and all subsequent ones.
Speculative decoding can accelerate reasoning and other test-time computations, but it can be overly strict in how it discards tokens (Bachmann et al., 2025; Pan et al., 2025; Tran-Thien, 2024). Intuitively, if a model generates a reasoning chain, not all mismatching tokens are equally important: errors in derivation should be fixed, while minor word choices should not. Judge Decoding (Bachmann et al., 2025) takes advantage of this by labeling which tokens are important for reasoning (and which are not) and allowing speculative decoding to accept more tokens by skipping the unimportant ones. However, their approach relies on human annotators to determine which tokens are important for reasoning. This complicates adoption and can be prone to human errors, particularly if the task requires expert knowledge (e.g., complex mathematical proofs or software engineering)
@@
Our experiments with Llama 3.x models demonstrate that the proposed approach can indeed identify important tokens and save time on speculation, accepting on average over 40 tokens per target model forward pass (approx. $2{\times}$ that of speculative decoding), at the cost of a ${\leq}1{\%}$ drop in accuracy on GSM8K (Cobbe et al., 2021) and even more with a minor accuracy drawdown. We obtain similar results with the Qwen2.5(Yang et al., 2024) family of models, observing comparable number of accepted tokens and accuracy trade-offs. When applied to programming tasks on LiveCodeBench (Jain et al., 2024), our approach is able to determine different task-specific important tokens, showing similar performance gains. The proposed framework is simple and general, using a classifier only when the original algorithm would reject a token, making it compatible with arbitrary speculative decoding algorithms. The main contributions of our work can be summarized as follows:
@@
  • We integrate AutoJudge with the vLLM framework (Kwon et al., 2023) and report the inference speed on A100 GPUs for both 8B and 70B target models, with up to $2{\times}$ speedup over speculative decoding at a ${\leq}1\%$ quality decrease, and on H100 GPUs with a 405B target model.
@@ 2 Background
Our work builds on top of speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023), a family of inference algorithms that accelerate token generation by improving hardware utilization. Speculative Decoding uses an auxiliary “draft” model to generate K>1K{>}1 possible future tokens, then runs the main “target” model in parallel to verify∗*∗**For greedy decoding, it checks that the drafted tokens are the same as the target model’s own next token predictions. For sampling, it uses a procedure that matches the sampling probabilities (Leviathan et al., 2023). the generated tokens. The drafted tokens that agree with the target model predictions are accepted by the algorithm. In turn, the first mismatching token and all subsequent ones are rejected. This way, the method guarantees that all generated tokens follow the same distribution as sampling from the target model. Subsequent works improve on this idea by generating draft trees instead of single sequences (Miao et al., 2023; Liu et al., 2023; Chen et al., 2024; Svirschevski et al., 2024), training specialized “heads” to draft next tokens based on the model’s hidden states (Cai et al., 2024; Ankner et al., 2024; Li et al., 2024d, c), and more (Fu et al., 2023; Spector and Re, 2023; Sun et al., 2023; He et al., 2023).
@@
The core guarantee of Speculative Decoding is that all generated tokens follow the probability distribution of the original model. However, there are practical scenarios where this guarantee can be sacrificed in favor of faster inference, which is known as lossy speculative decoding algorithms (Tran-Thien, 2024; Narasimhan et al., 2025; Kim et al., 2023). Our work extends one such method: Judge Decoding (Bachmann et al., 2025). The core idea of Judge Decoding is that speculative decoding should only reject the mismatching token if accepting it would harm the response quality. For instance, in mathematical reasoning, errors in the equations or logical fallacies are important for the final quality, while minor style changes are not. When writing code, algorithmic errors are important, while minor variable renames can be skipped in favor of faster inference.
The main challenge of Judge Decoding is determining which of the generated tokens can be skipped this way. Bachmann et al. (2025) address this problem by manually labeling a training dataset for the classifier. Judge Decoding requires human annotators to find the “mistake” — the first mismatching token that led the draft model to diverge from the original answer. The resulting dataset of high-quality training examples is then used to train a linear classifier that detects such “mistakes” during inference. Authors demonstrate that the collected dataset can, in principle, be reused across tasks and models. However, using the data from one task for inference on a different task results in substantial performance drawdown. Intuitively, different tasks (such as creative writing, math, or programming) have different criteria for which parts of the generated response matter most. Hence, it is best to train the important token classifier for the exact task at hand. However, doing so with Judge Decoding would require relabeling the data by human annotators, which can be costly and time-consuming in specialized domains such as medicine or law. To alleviate this problem, we develop an automated search procedure for determining important tokens without external human (or LLM) annotators.
@@ 3.1 Mining Important Tokens
Without the loss of generality, we also assume that there is a problem-specific way to extract the final answer from the model’s response, $a=\textsc{extractAnswer}(y)$. In mathematical reasoning tasks such as GSM8K (Cobbe et al., 2021), the final answer is literally whatever the model puts after "the final answer (is)". In programming tasks, the “answer” would be the output from the testing system given the generated code — either a report about passing and failing tests or a testing error (e.g., an Out Of Memory or Syntax Error). Finally, we say that two answers are equivalent $a_{ref}{\equiv}a_{alt}$ if they are the same from the downstream task perspective. Note that this does not require them to be exactly equal as strings: in math problems, $\texttt{1.5}{\equiv}\texttt{3/2}$, whereas in programming tasks, two programs can be equivalent despite having different variable names. If the task at hand does not have a formalized evaluation procedure, e.g., general conversation agents, we can define $\textsc{extractAnswer}(y)=y$ and detect if two answers are equivalent using an LLM or human judges.
@@
However, this approach has a significant downside in that it assumes that all subsequent tokens will be generated by $\theta_{target}$, whereas in reality, some of them may be generated by $\theta_{draft}$ following the same algorithm. In preliminary experiments, we found that, with a capable enough $\theta_{target}$, even significant generation errors can be detected and self-corrected (similar to the “Aha moment” from DeepSeek-AI et al. (2025); Muennighoff et al. (2025)). However, if the model makes multiple mistakes, they eventually reach a critical mass, leading to an incorrect answer.
@@ 3.3 Inference
Generality. In our initial experiments, we focus on traditional speculative decoding (Leviathan et al., 2023; Chen et al., 2023) for simplicity. However, our algorithm is compatible with arbitrary speculative decoding algorithms, including tree-based (Miao et al., 2023; Svirschevski et al., 2024; Chen et al., 2024) and single-model multi-head algorithms (Cai et al., 2024; Li et al., 2024d, c). This also means that our approach can be integrated into existing inference frameworks such as vLLM (Kwon et al., 2023), TensorRT-LLM (NVIDIA, 2023) or TGI (Hugging Face, 2023).
@@ 4 Experiments
We organize our evaluations as follows: in Section 4.1 we evaluate AutoJudge on the GSM8K (Cobbe et al., 2021) mathematical reasoning benchmark. Next, in Section 4.2, we evaluate our approach on programming tasks from LiveCodeBench (Jain et al., 2024). Finally, Section 4.3 contains GPU inference speed benchmarks with our vLLM implementation. We focus on two pairs of Llama 3.x models: 1) Llama-3.2-1B-Instruct draft / Llama-3.1-8B-Instruct target∗*∗**The reason why the two models have different minor versions (i.e. 3.1 and 3.2) is that the 3.2 version does not have the larger 8B models, and the 3.1 version does not have the smaller 1B models. and 2) Llama-3.1-8B-Instruct draft / Llama-3.1-70B-Instruct target. We also report results with Qwen2.5 models on the GSM8K benchmark to showcase the method’s transferability across model families. Our main experiments run in bfloat16 precision (see Appendix C). We run AutoJudge on top of the standard speculative decoding algorithm (Leviathan et al., 2023) in main experiments and explore EAGLE-2 in Appendix J.
@@ 4.1 Mathematical Reasoning with GSM8K
Our first set of experiments is based on the GSM8K dataset with grade school mathematical problems. This dataset has a natural split with ${\approx}7.47$K training samples and ${\approx}1.32$K test samples. Following the standard evaluation procedure, we use the training set to “mine” important tokens with Algorithm 1 and train the classifier, then run inference and evaluate on the test set with the recommended parameters (Gao et al., 2021) for zero-shot and 8-shot evaluation: greedy inference with a prompt that encourages chain-of-thought reasoning. During training, we consider two responses equivalent ($a{\equiv}\hat{a}$ in Algorithm˜1) if the extracted final answers (numbers) are equal. For reference, we provide an example of important token assignments found by our algorithm in Figure 2.
We train a classifier on the last hidden state embeddings from both draft and target models (concatenated) for encoded draft tokens. The training dataset contains ${\approx}130$K mismatches, about $20\%$ of which are deemed important. We train a logistic regression with the $L_{2}$ regularization coefficient tuned individually for each setup(for instance $10^{-4}$ for 8B/70B), found by grid search over a logarithmic grid between $10^{0},\dots,10^{-7}$. During inference, we integrate the trained classifier into the speculative decoding loop from Leviathan et al. (2023) during verification. Whenever the original algorithm would reject a token, we run the classifier to determine if changing that token affects the final response quality, and if not — accept the token and continue verification for subsequent tokens (if any). Since the resulting algorithm can accept additional tokens, we use the increased draft window size of W=64W{=}64 tokens for all evaluations.
@@
We report two main metrics: downstream accuracy and the number of accepted tokens per speculative decoding cycle. The accuracy is measured as the exact match rate for the final answer extracted from the response as per standard GSM8K evaluation protocol. In turn, we report the decoding speed in terms of the number of tokens accepted per target model forward pass, aiming to decouple our results from the specific hardware configuration. We evaluate AutoJudge with different classifier thresholds, balancing between accuracy and speed. Our baselines are traditional speculative decoding, decoding with the draft model, and a simpler lossy speculative decoding method. In the latter, we accept a mismatching draft token if it is within top-$K$ most likely tokens of the target model, similarly to how it is defined in Bachmann et al. (2025). We report $K{=}2,4,8,\dots,|V|$ for different speed-accuracy tradeoffs: increasing $K$ results in more accepted tokens but reduces accuracy.
@@ 4.2 Programming with LiveCodeBench
Next, we test if the AutoJudge search algorithm is able to generalize between domains. For this purpose, we evaluate the same model family on the LiveCodeBench (Jain et al., 2024) programming benchmark. Here, we use the code_generation_lite∗*∗**huggingface.co/datasets/livecodebench/code_generation_lite dataset with the version tag release_v5. The dataset contains 880 programming tasks; we evaluate on all three subsets: easy, medium and hard. Since LiveCodeBench does not have a dedicated training split, we evaluate using out-of-fold predictions. Namely, we split the dataset randomly into 5 folds. For each fold, we evaluate using the classifier trained on the 4 remaining folds. We use the standard evaluation protocol: extracting the generated code and evaluating it using the benchmark’s built-in test suite. We follow the standard evaluation protocol for this benchmark and report Pass@1 in zero-shot setting.
@@
The results in Figure 5 are similar to what we observed in Section 4.1: AutoJudge decoding can accept over 22 tokens per speculative decoding cycle at the cost of a ${\approx}2\%$ accuracy decrease. This results in approximately $2{\times}$ increase over traditional speculative decoding (Leviathan et al., 2023). The top-$K$ baseline can similarly achieve some increase in the number of accepted tokens, but AutoJudge decoding offers significantly better quality-speed tradeoffs across all configurations. We report additional configurations and threshold values in Appendix D.2, including the setup for AutoJudge decoding “out-of-domain”, i.e. using the classifier trained on GSM8K for LiveCodeBench evaluation. This out-of-domain configuration results in inferior performance, which aligns with our hypothesis that the important tokens depend on the problem type and evaluation criteria.

### v4

@@ 1 Introduction
Recent advances in LLM capabilities, including chain-of-thought reasoning (cot_wei_2022; zero_shot_cot_Kojima2022LargeLM; challenging_bigbench_solved_with_cot_Suzgun2022ChallengingBT), writing complex software (Rozire2023CodeLO; Li2023StarCoderMT; swe_survey_2024), or interacting with external tools (Schick2023ToolformerLM; Qin2023ToolLLMFL), increasingly rely on inference-time computation (scaling_test_time_snell2024scaling; beeching2024scalingtesttimecompute). This progress is further accelerated with the release of reasoning-capable models, both proprietary (openai_o1; AnthropicClaude3.7Sonnet; googledeepmind2025gemini25thinking) and open-access (deepseek_r1; meta2025llama4; qwq32b), that were explicitly trained to perform these kinds of inference-time computation. However, as LLMs tackle harder problems, they also tend to generate longer sequences (muennighoff2025s1) with tens of thousands of tokens (yeotong2025longcot), taking up tens of minutes (and hundreds of dollars in costs) per task (openai_arc_prize_o3).
A popular way to speed up LLM inference is through speculative decoding (leviathan2023fast; speculative_sampling_deepmind), which uses a small “draft” model to propose the likely next tokens, then verifies these tokens with the main model in parallel. This method, along with its successors (specinfer; cai2024medusa; li2024eagle), can speed up LLM inference while guaranteeing that the generated outputs match those of the original model (for greedy inference) or follow the same distribution. To achieve this, speculative decoding algorithms check if the draft tokens match the original model predictions. If there is a mismatch, they discard the incorrect token and all subsequent ones.
Speculative decoding can accelerate reasoning and other test-time computations, but it can be overly strict in how it discards tokens (bachmann2025judgedecoding; pan2025specreason; vivien2024optimal_lossy). Intuitively, if a model generates a reasoning chain, not all mismatching tokens are equally important: errors in derivation should be fixed, while minor word choices should not. Judge Decoding (bachmann2025judgedecoding) takes advantage of this by labeling which tokens are important for reasoning (and which are not) and allowing speculative decoding to accept more tokens by skipping the unimportant ones. However, their approach relies on human annotators to determine which tokens are important for reasoning. This complicates adoption and can be prone to human errors, particularly if the task requires expert knowledge (e.g., complex mathematical proofs or software engineering)
@@
Our experiments with Llama 3.x models demonstrate that the proposed approach can indeed identify important tokens and save time on speculation, accepting on average over 40 tokens per target model forward pass (approx. $2{\times}$ that of speculative decoding), at the cost of a ${\leq}1{\%}$ drop in accuracy on GSM8K (cobbe2021gsm8k) and even more with a minor accuracy drawdown. We obtain similar results with the Qwen2.5(qwen2.5) family of models, observing comparable number of accepted tokens and accuracy trade-offs. When applied to programming tasks on LiveCodeBench (jain2024livecodebenchholisticcontaminationfree), our approach is able to determine different task-specific important tokens, showing similar performance gains. The proposed framework is simple and general, using a classifier only when the original algorithm would reject a token, making it compatible with arbitrary speculative decoding algorithms. The main contributions of our work can be summarized as follows:
@@
  • We integrate AutoJudge with the vLLM framework (kwon2023efficient) and report the inference speed on A100 GPUs for both 8B and 70B target models, with up to $2{\times}$ speedup over speculative decoding at a ${\leq}1\%$ quality decrease, and on H100 GPUs with a 405B target model.
@@ 2 Background
Our work builds on top of speculative decoding (speculative_first; leviathan2023fast; speculative_sampling_deepmind), a family of inference algorithms that accelerate token generation by improving hardware utilization. Speculative Decoding uses an auxiliary “draft” model to generate K>1K{>}1 possible future tokens, then runs the main “target” model in parallel to verify∗*∗**For greedy decoding, it checks that the drafted tokens are the same as the target model’s own next token predictions. For sampling, it uses a procedure that matches the sampling probabilities (leviathan2023fast). the generated tokens. The drafted tokens that agree with the target model predictions are accepted by the algorithm. In turn, the first mismatching token and all subsequent ones are rejected. This way, the method guarantees that all generated tokens follow the same distribution as sampling from the target model. Subsequent works improve on this idea by generating draft trees instead of single sequences (specinfer; onlinespecdec; chen2024sequoiascalablerobusthardwareaware; specexec), training specialized “heads” to draft next tokens based on the model’s hidden states (cai2024medusa; ankner2024hydrasequentiallydependentdraftheads; li2024eagle; eagle2), and more (spec_lookahead; speculative_staged; spectr; spec_retrieval).
@@
The core guarantee of Speculative Decoding is that all generated tokens follow the probability distribution of the original model. However, there are practical scenarios where this guarantee can be sacrificed in favor of faster inference, which is known as lossy speculative decoding algorithms (vivien2024optimal_lossy; narasimhan2025fastercascades; biglittledecoder). Our work extends one such method: Judge Decoding (bachmann2025judgedecoding). The core idea of Judge Decoding is that speculative decoding should only reject the mismatching token if accepting it would harm the response quality. For instance, in mathematical reasoning, errors in the equations or logical fallacies are important for the final quality, while minor style changes are not. When writing code, algorithmic errors are important, while minor variable renames can be skipped in favor of faster inference.
The main challenge of Judge Decoding is determining which of the generated tokens can be skipped this way. bachmann2025judgedecoding address this problem by manually labeling a training dataset for the classifier. Judge Decoding requires human annotators to find the “mistake” — the first mismatching token that led the draft model to diverge from the original answer. The resulting dataset of high-quality training examples is then used to train a linear classifier that detects such “mistakes” during inference. Authors demonstrate that the collected dataset can, in principle, be reused across tasks and models. However, using the data from one task for inference on a different task results in substantial performance drawdown. Intuitively, different tasks (such as creative writing, math, or programming) have different criteria for which parts of the generated response matter most. Hence, it is best to train the important token classifier for the exact task at hand. However, doing so with Judge Decoding would require relabeling the data by human annotators, which can be costly and time-consuming in specialized domains such as medicine or law. To alleviate this problem, we develop an automated search procedure for determining important tokens without external human (or LLM) annotators.
@@ 3.1 Mining Important Tokens
Without the loss of generality, we also assume that there is a problem-specific way to extract the final answer from the model’s response, $a=\textsc{extractAnswer}(y)$. In mathematical reasoning tasks such as GSM8K (cobbe2021gsm8k), the final answer is literally whatever the model puts after "the final answer (is)". In programming tasks, the “answer” would be the output from the testing system given the generated code — either a report about passing and failing tests or a testing error (e.g., an Out Of Memory or Syntax Error). Finally, we say that two answers are equivalent $a_{ref}{\equiv}a_{alt}$ if they are the same from the downstream task perspective. Note that this does not require them to be exactly equal as strings: in math problems, $\texttt{1.5}{\equiv}\texttt{3/2}$, whereas in programming tasks, two programs can be equivalent despite having different variable names. If the task at hand does not have a formalized evaluation procedure, e.g., general conversation agents, we can define $\textsc{extractAnswer}(y)=y$ and detect if two answers are equivalent using an LLM or human judges.
@@
However, this approach has a significant downside in that it assumes that all subsequent tokens will be generated by $\theta_{target}$, whereas in reality, some of them may be generated by $\theta_{draft}$ following the same algorithm. In preliminary experiments, we found that, with a capable enough $\theta_{target}$, even significant generation errors can be detected and self-corrected (similar to the “Aha moment” from deepseek_r1; muennighoff2025s1). However, if the model makes multiple mistakes, they eventually reach a critical mass, leading to an incorrect answer.
@@ 3.3 Inference
Generality. In our initial experiments, we focus on traditional speculative decoding (leviathan2023fast; speculative_sampling_deepmind) for simplicity. However, our algorithm is compatible with arbitrary speculative decoding algorithms, including tree-based (specinfer; specexec; chen2024sequoiascalablerobusthardwareaware) and single-model multi-head algorithms (cai2024medusa; li2024eagle; eagle2). This also means that our approach can be integrated into existing inference frameworks such as vLLM (kwon2023efficient), TensorRT-LLM (trt-llm) or TGI (tgi).
@@ 4 Experiments
We organize our evaluations as follows: in Section 4.1 we evaluate AutoJudge on the GSM8K (cobbe2021gsm8k) mathematical reasoning benchmark. Next, in Section 4.2, we evaluate our approach on programming tasks from LiveCodeBench (jain2024livecodebenchholisticcontaminationfree). Finally, Section 4.3 contains GPU inference speed benchmarks with our vLLM implementation. We focus on two pairs of Llama 3.x models: 1) Llama-3.2-1B-Instruct draft / Llama-3.1-8B-Instruct target∗*∗**The reason why the two models have different minor versions (i.e. 3.1 and 3.2) is that the 3.2 version does not have the larger 8B models, and the 3.1 version does not have the smaller 1B models. and 2) Llama-3.1-8B-Instruct draft / Llama-3.1-70B-Instruct target. We also report results with Qwen2.5 models on the GSM8K benchmark to showcase the method’s transferability across model families. Our main experiments run in bfloat16 precision (see Appendix C). We run AutoJudge on top of the standard speculative decoding algorithm (leviathan2023fast) in main experiments and explore EAGLE-2 in Appendix J.
@@ 4.1 Mathematical Reasoning with GSM8K
Our first set of experiments is based on the GSM8K dataset with grade school mathematical problems. This dataset has a natural split with ${\approx}7.47$K training samples and ${\approx}1.32$K test samples. Following the standard evaluation procedure, we use the training set to “mine” important tokens with Algorithm 1 and train the classifier, then run inference and evaluate on the test set with the recommended parameters (eval-harness) for zero-shot and 8-shot evaluation: greedy inference with a prompt that encourages chain-of-thought reasoning. During training, we consider two responses equivalent ($a{\equiv}\hat{a}$ in Algorithm˜1) if the extracted final answers (numbers) are equal. For reference, we provide an example of important token assignments found by our algorithm in Figure 2.
We train a classifier on the last hidden state embeddings from both draft and target models (concatenated) for encoded draft tokens. The training dataset contains ${\approx}130$K mismatches, about $20\%$ of which are deemed important. We train a logistic regression with the $L_{2}$ regularization coefficient tuned individually for each setup(for instance $10^{-4}$ for 8B/70B), found by grid search over a logarithmic grid between $10^{0},\dots,10^{-7}$. During inference, we integrate the trained classifier into the speculative decoding loop from leviathan2023fast during verification. Whenever the original algorithm would reject a token, we run the classifier to determine if changing that token affects the final response quality, and if not — accept the token and continue verification for subsequent tokens (if any). Since the resulting algorithm can accept additional tokens, we use the increased draft window size of W=64W{=}64 tokens for all evaluations.
@@
We report two main metrics: downstream accuracy and the number of accepted tokens per speculative decoding cycle. The accuracy is measured as the exact match rate for the final answer extracted from the response as per standard GSM8K evaluation protocol. In turn, we report the decoding speed in terms of the number of tokens accepted per target model forward pass, aiming to decouple our results from the specific hardware configuration. We evaluate AutoJudge with different classifier thresholds, balancing between accuracy and speed. Our baselines are traditional speculative decoding, decoding with the draft model, and a simpler lossy speculative decoding method. In the latter, we accept a mismatching draft token if it is within top-$K$ most likely tokens of the target model, similarly to how it is defined in bachmann2025judgedecoding. We report $K{=}2,4,8,\dots,|V|$ for different speed-accuracy tradeoffs: increasing $K$ results in more accepted tokens but reduces accuracy.
@@ 4.2 Programming with LiveCodeBench
Next, we test if the AutoJudge search algorithm is able to generalize between domains. For this purpose, we evaluate the same model family on the LiveCodeBench (jain2024livecodebenchholisticcontaminationfree) programming benchmark. Here, we use the code_generation_lite∗*∗**huggingface.co/datasets/livecodebench/code_generation_lite dataset with the version tag release_v5. The dataset contains 880 programming tasks; we evaluate on all three subsets: easy, medium and hard. Since LiveCodeBench does not have a dedicated training split, we evaluate using out-of-fold predictions. Namely, we split the dataset randomly into 5 folds. For each fold, we evaluate using the classifier trained on the 4 remaining folds. We use the standard evaluation protocol: extracting the generated code and evaluating it using the benchmark’s built-in test suite. We follow the standard evaluation protocol for this benchmark and report Pass@1 in zero-shot setting.
@@
The results in Figure 5 are similar to what we observed in Section 4.1: AutoJudge decoding can accept over 22 tokens per speculative decoding cycle at the cost of a ${\approx}2\%$ accuracy decrease. This results in approximately $2{\times}$ increase over traditional speculative decoding (leviathan2023fast). The top-$K$ baseline can similarly achieve some increase in the number of accepted tokens, but AutoJudge decoding offers significantly better quality-speed tradeoffs across all configurations. We report additional configurations and threshold values in Appendix D.2, including the setup for AutoJudge decoding “out-of-domain”, i.e. using the classifier trained on GSM8K for LiveCodeBench evaluation. This out-of-domain configuration results in inferior performance, which aligns with our hypothesis that the important tokens depend on the problem type and evaluation criteria.